A population consisting of 6 measurements 4 ̧ 5 ̧ 8 ̧ 12 ̧ 9 and 12.

a. Solve for the mean of the population.

b. Solve for the variance of the population.

c. Suppose samples of size 4 are to be taken from this population ̧ what is

the mean of the sampling distribution?

d. What is its variance of the sampling distribution?

e. What is the standard error of the mean?

f. Determine whether the estimate for the population mean is good or bad.

Support your answer.


1
Expert's answer
2022-05-04T11:49:38-0400

We have population values 4, 5, 8, 12, 9, 12, population size N=6 and sample size n=4.

a. Mean of population "(\mu)" = "\dfrac{4+5+8+12+9+12}{6}=25/3"

b. Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{N}=\dfrac{86}{9}"


"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{86}{9}}=\dfrac{\sqrt{86}}{3}"


c. The number of possible samples which can be drawn without replacement is "^{N}C_n=^{6}C_4=15."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 4,5,8,12 & 29/4 \\
 \hdashline
 2 & 4,5,8,9 & 26/4 \\
 \hdashline
 3 & 4,5,8,12 & 29/4 \\
 \hdashline
 4 & 4,5,12,9 & 30/4 \\
 \hdashline
 5 & 4,5,12,12 & 33/4 \\
 \hdashline
 6 & 4,5,9,12 & 30/4 \\
 \hdashline
 7 & 4,8,12,9 & 33/4 \\
 \hdashline
 8 & 4,8,12,12 & 36/4 \\
 \hdashline
 9 & 4,8,9,12 & 33/4 \\
 \hdashline
 10 & 4,12,9,12 & 37/4 \\
 \hdashline
 11 & 5,8,12,9 & 34/4 \\
 \hdashline
 12 & 5,8,12,12 & 37/4 \\
 \hdashline
 13 & 5,8,9,12 & 34/4 \\
 \hdashline
 14 & 5,12,9,12 & 38/4 \\
 \hdashline
 15 & 8,12,9,12 & 41/4 \\
 \hdashline
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) &\bar{X}^2 f(\bar{X}) 
\\ \hline
 26/4 & 1/15 & 26/60 & 676/240 \\
 \hdashline
 29/4 & 2/15 & 58/60 & 1682/240 \\
 \hdashline
 30/4 & 2/15 & 60/60 & 1800/240 \\
 \hdashline
 33/4 & 3/15 & 99/60 & 3267/240 \\
 \hdashline
 34/4 & 2/15 & 68/60 & 2312/240 \\
 \hdashline
 36/4 & 1/15 & 36/60 & 1296/240 \\
 \hdashline
 37/4 & 2/15 & 74/60 & 2738/240 \\
 \hdashline
 38/4 & 1/15 & 38/60 & 1444/240 \\
 \hdashline
 41/4 & 1/15 & 41/60 & 1681/240 \\
 \hdashline
\end{array}"


Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=25/3"


d. The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{16896}{240}-(\dfrac{25}{3})^2=\dfrac{43}{45}"

e.

"\sigma_{\bar{X}}=\sqrt{\dfrac{43}{45}}\approx0.9775"

f.


"\mu_{\bar{X}}=E(\bar{X})=25/3=\mu"

"Var(\bar{X})=\sigma^2_{\bar{X}}=\dfrac{43}{45}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"